OpenStudy (anonymous):

Find the probability, if you are dealt two cards successively (with replacement of the first) from a standard 52-card deck, find the probability of getting a heart on the first card and a diamond on the second.

OpenStudy (anonymous):

52 cards 13 hearts, 13 diamonds, 13 spades, 13 clubs P(Heart) = 13/52. You replace the card P(Diamond) = 13/52. Independent events, multiply the probabilities. P(Heart then Diamond) = 13/52*13/52 = 1/4*1/4 = 1/16 Is this the right answer?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

Thnx :)